# A deep solver for backward stochastic Volterra integral equations

### v3

@@ Abstract
We present the first deep-learning solver for backward stochastic Volterra integral equations (BSVIEs) and their fully-coupled forward-backward variants. The method trains a neural network to approximate the two solution fields in a single stage, avoiding the use of nested time-stepping cycles that limit classical algorithms. For the decoupled case we prove a non-asymptotic error bound composed of an a posteriori residual plus the familiar square root dependence on the time step. Numerical experiments confirm this rate and reveal two key properties: scalability, in the sense that accuracy remains stable from low dimension up to 500 spatial variables while GPU batching keeps wall-clock time nearly constant; and generality, since the same method handles coupled systems whose forward dynamics depend on the backward solution. These results open practical access to a family of high-dimensional, time-inconsistent problems in stochastic control and quantitative finance.
@@ 3.2 A time discrete variational formulation
We assume that the expectation can be approximated to any desired accuracy. In practice we use Monte Carlo simulation, whose mean–square error decreases at the classical rate $O\!\left(M^{-1}\right)$ when the sample size $M$ grows.
@@ 6 Numerical experiments
Although we report results with a single feed-forward architecture for all problems, during development we experimented with a wide range of hyper-parameters: between $1$ and $6$ hidden layers, 1010300300 neurons per layer, ReLU versus $\tanh$ activations, and both Adam and SGD optimizers. Once the network had sufficient capacity, the solver’s accuracy and convergence rate were essentially insensitive to the exact depth, width, or activation choice. We therefore settled on a network with three hidden layers and $100$ neurons per layer, oversized for the low-dimensional examples yet still tractable in up to 500 dimensions, which performed on par with both slimmer and deeper alternatives. This configuration is thus reported as the smallest architecture that remained robust across all test cases.
@@ 6.1.3 Example 1: Empirical error analysis
$\begin{split}\mathrm{Err}^{Y}(N,M)&\vcentcolon=\frac{1}{M}\sum_{m=1}^{M}\sum_{n=0}^{N-1}\Big|Y_{t_{n}}(m)-Y_{n}^{\pi,\theta^{*}}(m)\Big|^{2}h,\\ \mathrm{Err}^{Z}(N,M)&\vcentcolon=\frac{1}{M}\sum_{m=1}^{M}\sum_{k=0}^{N-1}\sum_{n=k}^{N-1}\Big|Z_{t_{k},t_{n}}(m)-Z_{k,n}^{\pi,\theta^{*}}(m)\Big|^{2}h^{2},\\ \mathrm{Err}^{\text{T}}(N,M)&\vcentcolon=\frac{1}{M}\sum_{m=1}^{M}\sum_{n=0}^{N-1}\Big|\mathscr{Y}_{N}^{\pi,\theta^{*}}(n)(m)-g\big(t_{n},X_{n}^{\pi,\theta^{*}}(m),X_{N}^{\pi,\theta^{*}}(m)\big)\Big|^{2}.\end{split}$ (24)
@@
In both Example 1A and Example 1B, we observe an empirical convergence order of 1 for the optimization loss. However, for $Y$ and $Z$, the convergence order is 1 in Example 1A and 0.5 in Example 1B. This suggests that employing a higher-order approximation for the FSDE can improve the overall convergence order achieved for the BSVIE.
@@ Acknowledgments
$\mathbb{E}\!\left|\delta Y_{n}^{k}-\delta f_{n}^{k}h\right|^{2}$ $\geq\mathbb{E}|\delta Y_{n}^{k}|^{2}-2h\,\mathbb{E}[\delta Y_{n}^{k}\,\delta f_{n}^{k}]$

### v4

@@ Abstract
We present the first deep-learning solver for backward stochastic Volterra integral equations (BSVIEs) and their fully-coupled forward-backward variants. The method trains a neural network to approximate the two solution fields in a single stage, avoiding the use of nested time-stepping cycles that limit classical algorithms. For the decoupled case we prove a non-asymptotic error bound composed of an a posteriori residual plus the familiar square root dependence on the time step. Numerical experiments are consistent with this rate and reveal two key properties: scalability, in the sense that accuracy remains stable from low dimension up to 500 spatial variables while GPU batching keeps wall-clock time nearly constant; and generality, since the same method handles coupled systems whose forward dynamics depend on the backward solution. These results open practical access to a family of high-dimensional, time-inconsistent problems in stochastic control and quantitative finance.
@@ 3.2 A time discrete variational formulation
Expectation.
We approximate the expectations appearing in the loss by Monte Carlo simulation. For any fixed parameter choice, the Monte Carlo estimator of the expected loss is unbiased, with mean–square error $O(M^{-1})$ (equivalently, standard error $O(M^{-1/2})$) as the number of samples $M$ increases. Note that this rate concerns the accuracy of the Monte Carlo approximation of the expectation operator, not the convergence of the loss itself. The loss function is already a mean–squared quantity by definition, what converges with rate $O(M^{-1})$ is the estimator of that expectation, not the loss or its optimizer.
@@ 6 Numerical experiments
Although we report results with a single feed-forward architecture for all problems, during development we experimented with a wide range of hyper-parameters: between $1$ and $6$ hidden layers, 1010300300 neurons per layer, ReLU versus $\tanh$ activations, and both Adam and SGD optimizers. Once the network had sufficient capacity, the solver’s accuracy and empirical convergence rate were essentially insensitive to the exact depth, width, or activation choice. We therefore settled on a network with three hidden layers and $100$ neurons per layer, oversized for the low-dimensional examples yet still tractable in up to 500 dimensions, which performed on par with both slimmer and deeper alternatives. This configuration is thus reported as the smallest architecture that remained robust across all test cases.
@@ 6.1.3 Example 1: Empirical error analysis
$\begin{split}\mathrm{Err}^{Y}(N,M)&\vcentcolon=\frac{1}{M}\sum_{m=1}^{M}\sum_{n=0}^{N-1}\Big|Y_{t_{n}}(m)-Y_{n}^{\pi,\theta^{*}}(m)\Big|^{2}h,\\ \mathrm{Err}^{Z}(N,M)&\vcentcolon=\frac{1}{M}\sum_{m=1}^{M}\sum_{k=0}^{N-1}\sum_{n=k}^{N-1}\Big|Z_{t_{k},t_{n}}(m)-Z_{k,n}^{\pi,\theta^{*}}(m)\Big|^{2}h^{2},\\ \mathrm{Err}^{\text{T}}(N,M)&\vcentcolon=\frac{1}{M}\sum_{m=1}^{M}\sum_{n=0}^{N-1}\Big|\mathscr{Y}_{N}^{\pi,\theta^{*}}(n)(m)-g\big(t_{n},X_{n}^{\pi,\theta^{*}}(m),X_{N}^{\pi,\theta^{*}}(m)\big)\Big|^{2}h.\end{split}$ (24)
@@
In both Example 1A and Example 1B, we observe an empirical convergence order of 1 for the optimization loss. However, for $Y$ and $Z$, the convergence order is 1 in Example 1A and 0.5 in Example 1B. This suggests that employing a higher-order approximation for the FSDE can improve the observed overall error rate.
@@ Acknowledgments
$\mathbb{E}\big|\delta Y_{n}^{k}-\delta f_{n}^{k}h\big|^{2}$ $\geq\mathbb{E}|\delta Y_{n}^{k}|^{2}-2h\,\mathbb{E}[\delta Y_{n}^{k}\,\delta f_{n}^{k}]$